OpenStudy (anonymous):

if the ratio of a circle's sector to its total area is 7/16,what is the measure of its sector's arc? A.181.8 B.205.3 C.157.5 D.none of the above

OpenStudy (anonymous):

circle's sector/ total area= 7/16 so, circle's sector=7/16 times total area

OpenStudy (anonymous):

i think total area is also given in this question , can you check it?

OpenStudy (anonymous):

total area is not given

OpenStudy (anonymous):

here, we take total area as 360 so, circle's sector=7/16 times 360

OpenStudy (anonymous):

so 157.5

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

:)